# A LoD of Gaussians: Unified Training and Rendering for Ultra-Large Scale Reconstruction with External Memory

### v3

@@ 4.1 Sequential Point Trees for Gaussian Splatting
Figure 1: An SPT and Gaussian hierarchy represent the same 5 Gaussians at varying levels of detail shown with four possible hierarchy cuts in red. Vertical lines in the SPT show the binary search result and horizontal lines the distance cut.
@@
Sequential Point Trees (SPT) were originally developed for point cloud LoD rendering, but can be trivially extended to ellipsoids and Gaussians. They enforce a more constrained but more efficient cut condition than Hierarchical:
@@
$c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$ (4)
@@
This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$. It requires storing only sorted pairs $\big(m_{d}(i),m_{d}(\text{parent}(i))\big)$, significantly reducing memory compared to full Gaussian hierarchies. To optimize cuts, Gaussians are sorted by $m_{d}(\text{parent}(i))$ in descending order. A binary search determines the cutoff index $N$, above which Gaussians are too fine to be rendered. Note that cuts are guaranteed to be proper, with nodes where $m_{d}(\text{parent}(i))>m_{d}(i)$ never being selected for the cut. Evidently, the level of detail of all Gaussians in the SPT is dictated by the camera’s distance to its root node. This can lead to Gaussians with $m_{d}(i)>\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view. To counteract this issue, we define: $M_{d}(i)=m_{d}(i)+\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$ as a conservative minimum distance function. By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\bm{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\bm{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$. In turn, this means that Gaussians that are further away from the camera than the root node will be selected at a higher level of detail. SPTs are best suited for tightly grouped Gaussians observed from distances greater than their mutual spacing. Their compact memory footprint and parallel evaluation make them well-suited for large-scale scenes. Figure 1 visualizes both hierarchy types and their LoD cuts for a toy example.
4.2 Densification
Figure 2: Example of densifying and respawning leaf nodes.
@@ 4.2 Densification
We take inspiration from 3DGS-MCMC (3DGSasMCMC), which ‘splits’ Gaussians, replacing them with two new Gaussians which together should appear similarly to the original. Notably, this mirrors how a parent node in a Gaussian hierarchy approximates its children. Therefore, we adopt this approach and instead ‘spawn’ two new child nodes for a Gaussian, increasing the size of the hierarchy with minimal artifacts.
Instead of pruning, 3DGSasMCMC declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian. We propose a similar strategy: when a leaf node dies, its parent is replaced by its sibling node; the dead leaf node and its parent are then respawned as children to another node, which is selected to be densified. See Figure 2 for an overview of the two hierarchy densification operations. Together, they ensure that the hierarchy can be expanded during training in a stable manner and rebalanced as required.
While 3DGSasMCMC choose Gaussians to densify using a random selection weighted by opacity, we employ the strategy from Hierarchical, which selects Gaussians for densification based on their maximal screen-space gradient. Avoiding destructive opacity resets, as advocated by 3DGS, is particularly important in our setting, as many Gaussians will no longer conform to the hierarchical structure after opacity is restored to normal. Further details on our densification method can be found in Appendix A.7.
4.3 The hierarchical SPT datastructure
This section covers previous approaches to LoD selection and why a new datastructure—the hierarchical SPT—is necessary for robust and efficient training.
@@ 4.3 The hierarchical SPT datastructure
Computing the cut set of a large Gaussian hierarchy is a costly operation that must be performed for every frame. A straightforward solution is to run a breadth-first search (BFS) from the root. This guarantees a proper cut and enables early pruning of large subtrees, e.g., via frustum culling. However, graph traversal is not well suited to parallel execution on the GPU.
@@
To enable GPU-accelerated cuts, Hierarchical evaluate the cut condition in parallel for each Gaussian:
$\Big(m_{d}(i)<\left\|\bm{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big)\land\Big(m_{d}(parent(i))\geq\left\|\bm{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big),$ (5)
@@
Figure 3: A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
We present the hierarchical SPT data structure, which combines the benefits of both approaches. To construct it from a Gaussian hierarchy, we cut it using a BFS on the condition $c_{\text{HSPT}}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$ with volume threshold size. The resulting cut set $\mathbb{C}_{\text{HSPT}}$ partitions the hierarchy into the upper hierarchy, which includes all Gaussians with volume greater than size, and the lower hierarchy, consisting of the subtrees rooted at the nodes in the cut set.
The volume of each root node in the lower hierarchy is now bounded by size, which also roughly bounds the extent of all Gaussians in the subtree. This provides an upper bound on the error introduced if the subtree is converted into an SPT. Consequently, each subtree of sufficient size in $\mathbb{C}_{\text{HSPT}}$ can be transformed into a Sequential Point Tree to accelerate cut computation. The HSPT-based cutting process then proceeds in two steps: first, a BFS on the upper hierarchy selects the required nodes and leaf/SPT subtrees for the current view. Second, each selected SPT is cut according to the camera’s distance to its root node. Together, these yield the full set of Gaussians needed for rendering the current frame. The construction and cutting process of an HSPT is illustrated in Figure 3. Figure 10 and Figure 15 show example frames with respectively SPTs highlighted and different levels of detail.
Rebuilding the HSPT for every training iteration would eliminate any performance benefit. Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus only needs to be updated infrequently. In practice, we rebuild the HSPT only after each densification step.
This infrequent recomputation allows us to use a more accurate—albeit more expensive—minimum distance metric than the inverse of maximal scale. Specifically, we define:
$m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$ (6)
@@
The major benefits of using BFS to cut the upper hierarchy are the guarantee of a proper cut and early culling of subtrees. To this end, we frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius equal to $(3\cdot\max_{j}s_{i}^{j})$ intersects with the view frustum. Using the Gaussian scale as a proxy for the extent of its entire subtree is not perfectly accurate, but comparisons to using a full bounding sphere hierarchy showed no discernable difference in our experiments. It should be noted that Gaussians are implicitly frustum culled during rasterization, but this early culling accelerates the cutting procedure and significantly decreases the number of Gaussians that need to be loaded from RAM, as can be seen in Figure 9.
4.4 Caching on the GPU
Figure 4: Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits. After optimization, newly accessed SPTs are added to the GPU cache.
@@ 4.4 Caching on the GPU
Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$-th SPT. During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\bm{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is cached, using a simple distance ratio tolerance:
$D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$ (7)
Here, $D_{\text{max}}$ defines the allowable range for using coarser-than-optimal detail, while $D_{\text{min}}$ limits how much finer detail can be tolerated. If this condition is met, the cached SPT cut is reused, avoiding a costly RAM-to-GPU transfer. While this heuristic introduces slight variability in rendered detail—since the LoD may depend on the cache state—we find that this stochasticity actually improves training robustness. In particular, the subtle variation in detail across views discourages overfitting to fixed camera distances and promotes better generalization across scales.
@@
To this end, we precompute a $k$-nearest-neighbour graph over all training view positions, where edge weights $w_{ij}$ correspond to the Euclidean distance between views $i$ and $j$. The next training view $j$ is then sampled from the $k$-nearest neighbours of the current view $i$ according to the distribution: $\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$ where $W$ is a normalization constant that also controls the degree of exploration.
However, care must be taken when modifying the order of training views, as deviating from uniform random sampling may introduce training bias. To counteract this, we inject a randomly selected view every 128 iterations, which we find sufficient to preserve generalization performance.
4.5 Memory Layout
Figure 5: Peak memory consumption of CPU and GPU for a training iteration on MC-smaller-city+ with 60 million Gaussians.
Figure 5 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MC-smaller-city+ dataset. The output frame from this iteration is shown in Figure 13. The majority of RAM usage is consumed by per-Gaussian properties and their corresponding ADAM optimizer states. In contrast, the hierarchy structure itself accounts for less than 10% of the total RAM footprint. On the GPU, the SPT metadata for all 60 million Gaussians occupies just 680 MB of VRAM and the upper hierarchy negligible 24 MB.
@@ 4.5 Memory Layout
The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as general PyTorch overhead. Figure A.1 provides a breakdown of all data associated with a single Gaussian and how they are distributed across CPU and GPU.
5 Evaluation
Our method is designed to enable seamless training and rendering on ultra-large-scale scenes comprising tens of thousands of views captured at vastly different scales. Unfortunately, most datasets of sufficient size contain either street-level or aerial views—but not both. To overcome these challenges, we turn to the synthetic MatrixCity dataset (MatrixCityDataset), which provides dense multi-scale views along with ground-truth camera poses and sparse geometry. We also include results on the real-world street-level datasets from H-3DGS. As baseline, we compare to recent large-scale 3DGS methods CityGaussian (CityGaussian), H-3DGS (Hierarchical), HorizonGS (HorizonGS) and OctreeGS (OctreeGS). We note that HorizonGS is not evaluated on the H-3DGS dataset, since it requires aerial views for training. Further results and additional details on the datasets and baselines and are included in A.4. Evaluations were run on a single H200 GPU with 141 GB VRAM, while all of our models were trained on a single RTX 3090 graphics card with 24 GB VRAM. The contribution of our components is ablated in Appendix A.3.
5.1 Results
Qualitative comparisons are shown in Figure 6 and quantitative comparisons in Tables 1 and 2. The outputs of H-3DGS and HorizonGS show significant floating and ghosting artifacts. While H-3DGS performs well on individual chunks of MC-smaller-city+, oversized floaters survive the merging procedure and obscure the majority of test images (cf. Figure 11, 13 and 17). The merging procedure of CityGaussian (CityGaussian), which is designed for aerial-only datasets, discards most of the trained Gaussians to avoid chunk artifacts, leading to mostly artifact-free views, but significantly blurry results. A LoD of Gaussians reconstructs novel views with a high degree of detail while avoiding chunk-based artifacts (cf. Figure 16), resulting in a considerable lead in quality metrics. Seamless training enables faster convergence during training, reducing the required number of training iterations considerably over divide-and-conquer based methods.
Table 1: MatrixCity novel view synthesis results. Results with * use the MC-smaller-city+ dataset due to memory constraints. Results with require suboptimal COLMAP initialization instead of the dataset’s provided point cloud. VRAM usage is measured while rendering the test images. Methods are considered out-of-memory (OOM) if they exceed 141GB of VRAM during rendering or training.
MC-small-city+ (42.2k images) / *MC-smaller-city+ (15.1k images)
@@ 5.1 Results
CityGaussian 19.78 0.650 0.475 2.4 GB 1.11M 2.7M
HorizonGS 12.06 0.521 0.544 20.6 GB 2.5M 50.3M
H-3DGS OOM OOM OOM >>141GB OOM OOM
OctreeGS OOM OOM OOM >>141GB OOM OOM
Ours 21.59 0.711 0.396 25.0GB 600k 136.7M
HorizonGS* 15.28 0.601 0.435 13.91 1.2M 34.1M
OctreeGS* OOM OOM OOM >>141GB OOM OOM
Ours* 21.71 0.710 0.370 17.7GB 310k 60M
H-3DGS* 14.57 0.516 0.563 97GB 4M 83.1M
Ours* 20.59 0.658 0.468 21.7GB 310k 24.7M
Table 2: Hierarchical 3DGS dataset novel view synthesis results.
Small City (5.8k images) Campus (22.0k images)

### v4

@@ 4.1. Sequential Point Trees for Gaussian Splatting
Figure 2. Visualization of Hierarchy and Densification. (2(a)) An SPT and Gaussian hierarchy show the same 5 Gaussians at varying levels of detail (red lines indicate cuts). Vertical lines show binary search results; horizontal lines show distance cuts. (2(b)) Leaf node densification and respawning.
@@
Sequential Point Trees (Dachsbacher et al., 2003) were originally developed for point cloud LoD rendering, but can be trivially extended to ellipsoids and Gaussians. They enforce a more constrained but more efficient cut condition than Kerbl et al. (2024):
@@
(4) $\scriptsize c_{\text{SPT}}(i,\text{cam})=m_{d}(\text{parent}(i))>\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}\geq m_{d}(i).$
@@
This condition is evaluated for all Gaussians in parallel, using the shared root-camera distance $\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$. It requires storing only sorted pairs $\big(m_{d}(i),m_{d}(\text{parent}(i))\big)$, significantly reducing memory compared to full Gaussian hierarchies. To optimize cuts, Gaussians are sorted by $m_{d}(\text{parent}(i))$ in descending order. A binary search determines the cutoff index $N$, above which Gaussians are too fine to be rendered. Note that cuts are guaranteed to be proper, with nodes where $m_{d}(\text{parent}(i))>m_{d}(i)$ never being selected for the cut. Evidently, the level of detail of all Gaussians in the SPT is dictated by the camera’s distance to its root node. This can lead to Gaussians with $m_{d}(i)>\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}$ being rendered, even though they would be too coarse for the current view. To counteract this issue, we define: $M_{d}(i)=m_{d}(i)+\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2},$ as a conservative minimum distance function. By the triangle inequality, selecting Gaussians satisfying $M_{d}(i)\leq\left\|\boldsymbol{\mu}_{\text{root}}-\mathbf{p}_{\text{cam}}\right\|_{2}$ guarantees $m_{d}(i)\leq\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{\text{cam}}\right\|_{2}$. In turn, this means that Gaussians that are further away from the camera than the root node will be selected at a higher level of detail. SPTs are best suited for tightly grouped Gaussians observed from distances greater than their mutual spacing. Their compact memory footprint and parallel evaluation make them well-suited for large-scale scenes. Figure 2 visualizes both hierarchy types and their LoD cuts.
4.2. Densification
@@
We take inspiration from 3DGS-MCMC (Kheradmand et al., 2024), which ‘splits’ Gaussians, replacing them with two new Gaussians which together should appear similarly to the original. Notably, this mirrors how a parent node in a Gaussian hierarchy approximates its children. Motivated by this correspondence, we adopt this approach and instead ‘spawn’ two new child nodes for a Gaussian, increasing the size of the hierarchy with minimal artifacts. This procedure leads to gradual increase in detail during densification, thereby avoiding the instability on large scenes observed in (Kerbl et al., 2024).
Instead of pruning, 3DGS-MCMC declares Gaussians below a certain opacity threshold as ‘dead’, and respawns them at the position of a high-opacity Gaussian. We propose a similar strategy: when a leaf node dies, its parent is replaced by its sibling node; the dead leaf node and its parent are then respawned as children to another node, which is selected to be densified. See Figure 2(b) for an overview of the two hierarchy densification operations. Together, they ensure that the hierarchy can be expanded during training in a stable and valid manner while being rebalanced as required.
While 3DGS-MCMC choose Gaussians to densify using a random selection weighted by opacity, we employ the strategy from Kerbl et al. (2024), which selects Gaussians for densification based on their maximal screen-space gradient. This criterion better aligns densification with view-dependent reconstruction error in large-scale scenes. For further details on densification, see Suppl. A.6.
4.3. The Hierarchical SPT Datastructure
We first review previous LoD selection approaches and motivate the need for a new datastructure—the hierarchical SPT—for robust and efficient training.
@@ 4.3. The Hierarchical SPT Datastructure
Computing the cut set of a large Gaussian hierarchy is costly and must be done for every frame. A straightforward solution is a BFS from the root, which guarantees a proper cut and enables early pruning of large subtrees (e.g., via frustum culling). However, graph traversal poorly suited to parallel execution on the GPU, making this approach prohibitively expensive at scale.
@@
To enable GPU-accelerated cuts, Kerbl et al. (2024) evaluate the cut condition in parallel for each Gaussian:
$\Big(m_{d}(i)<\left\|\boldsymbol{\mu}_{i}-\mathbf{p}_{cam}\right\|_{2}\Big)\land\Big(m_{d}(\text{parent}(i))\geq\left\|\boldsymbol{\mu}_{\text{parent}(i)}-\mathbf{p}_{cam}\right\|_{2}\Big)$,
@@
Figure 3. A Gaussian hierarchy is converted to an HSPT by cutting according to Gaussian volume and converting sufficiently large subtrees to SPTs. The HSPT can then be cut in a 2-step process.
Our HSPT data structure combines the benefits of both approaches. To construct it from a Gaussian hierarchy, we cut it using a BFS on the condition $c_{\text{HSPT}}(i)=s_{i}^{1}\cdot s_{i}^{2}\cdot s_{i}^{3}<\texttt{size}$ with volume threshold size. The resulting cut set $\mathbb{C}_{\text{HSPT}}$ partitions the hierarchy into the upper hierarchy, which includes all Gaussians with volume greater than size, and the lower hierarchy, consisting of the subtrees rooted at the nodes in the cut set.
The volume of each root node in the lower hierarchy is now bounded by size, which also roughly bounds the extent of all Gaussians in the subtree. This provides an upper bound on the error introduced if the subtree is converted into an SPT. Consequently, each subtree of sufficient size in $\mathbb{C}_{\text{HSPT}}$ can be transformed into an SPT to accelerate cut computation without violating cut correctness at higher levels. The HSPT-based cutting process then proceeds in two steps: first, a BFS on the upper hierarchy selects the required nodes and leaf/SPT subtrees for the current view. Second, each selected SPT is cut according to the camera’s distance to its root node. Together, these yield the full set of Gaussians needed for rendering the current frame. The construction and cutting process of an HSPT is illustrated in Figure 3. Figure 11 and 14 show example frames with respectively SPTs highlighted and different levels of detail.
Rebuilding the HSPT every training iteration would eliminate its performance benefits. Instead, we exploit the fact that the minimum distance $m_{d}$ evolves slowly during optimization and thus requires only infrequent updates. In practice, we rebuild the HSPT only after each densification step. This infrequent recomputation allows us to use a more accurate—albeit more expensive—minimum distance metric than the inverse of maximal scale. Specifically, we define:
(6) $m_{d}^{\prime}(i)=\frac{T}{\sqrt{s_{i}^{1}\cdot s_{i}^{2}+s_{i}^{1}\cdot s_{i}^{3}+s_{i}^{2}\cdot s_{i}^{3}}},$
@@
The main benefits of using BFS to cut the upper hierarchy are the guarantee of a proper cut and early culling of subtrees. We therefore frustum cull every node considered in the BFS by checking if a sphere around the Gaussians with radius $3\cdot\max_{j}s_{i}^{j}$ intersects the view frustum. While using the Gaussian scale as a conservative proxy for the entire subtree extent is not perfectly accurate, we observed no discernable difference compared to a full bounding sphere hierarchy in our experiments. Although Gaussians are implicitly frustum culled during rasterization, this early culling accelerates cut computation and significantly reduces the number of Gaussians loaded from RAM (see Figure 10).
4.4. Caching on the GPU
@@
Rather than caching individual Gaussians, we store the Gaussians from SPT cuts along with the cached distance from the camera to the root of each SPT, denoted $\bar{d}^{j}$ for the $j$-th SPT. During rendering, when the upper hierarchy is cut and the required SPTs identified, we compute $d^{j}=\|\boldsymbol{\mu}_{\text{root}(j)}-\mathbf{p}_{\text{cam}}\|_{2}$ and check whether a matching cut is cached, using a simple distance ratio tolerance:
Figure 4. Gaussians required for the current training view are assembled from three sources: the upper tree, newly loaded SPT cuts from RAM, and cache hits. After optimization, newly accessed SPTs are added to the GPU cache.
(7) $D_{\text{min}}\leq\frac{d^{j}}{\bar{d}^{j}}\leq D_{\text{max}}.$
Here, $D_{\text{max}}$ defines the allowable range for using coarser-than-optimal detail, while $D_{\text{min}}$ limits how much finer detail can be tolerated. If this condition is met, the cached SPT cut is reused, avoiding a costly RAM-to-GPU transfer.
While this heuristic introduces slight variability in rendered detail—since the LoD may depend on the cache state—we find that this stochasticity actually improves training robustness. In particular, subtle variations in detail across views discourage overfitting to fixed camera distances and promotes generalization across scales.
@@
To this end, we precompute a $k$-nearest-neighbour graph over all training view positions, where edge weights $w_{ij}$ correspond to the Euclidean distance between views $i$ and $j$. The next training view $j$ is then sampled from the $k$-nearest neighbours of the current view $i$ according to the distribution: $\mathbb{P}(j\mid i)\propto\frac{1}{w_{ij}+W},$ where $W$ is a normalization constant that also controls the degree of exploration. However, care must be taken when deviating from a uniformly sampled training view selection as this may introduce bias. To counteract this, we inject a randomly selected view every 128 iterations, which we find sufficient to preserve generalization performance.
4.5. Memory Layout
Figure 5. Peak memory consumption of CPU and GPU for a training iteration on MC-smaller-city+ with 60 million Gaussians.
Figure 5 illustrates the peak memory usage for a single training iteration of a 60-million-Gaussian hierarchy on the MC-smaller-city+ dataset. The majority of RAM usage is consumed by per-Gaussian properties and their corresponding ADAM optimizer states. In contrast, the hierarchy structure itself accounts for less than 10% of the total RAM footprint. On the GPU, the SPT metadata for all 60 million Gaussians occupies just 680 MB of VRAM and the upper hierarchy negligible 24 MB.
@@ 4.5. Memory Layout
The remaining GPU memory usage consists of auxiliary data, including cache management, hierarchy cut tracking, training and ground-truth images, as well as general PyTorch overhead. Figure 9 provides a breakdown of all data associated with a single Gaussian and how they are distributed across CPU and GPU.
Table 1. Novel view synthesis results. Results with require suboptimal COLMAP initialization instead of the provided point cloud. VRAM usage is measured while rendering the test images. Methods are considered out-of-memory (OOM) if they exceed 141GB of VRAM during training or rendering.
@@
MC-small-city+ (42.2k images)
CityGaussian 19.78 0.650 0.475 2.4GB 4.1GB 1.11M 2.7M
HorizonGS 12.06 0.521 0.544 20.6GB 25.3GB 2.5M N/A
OctreeGS OOM OOM OOM OOM OOM OOM OOM
H-3DGS OOM OOM OOM OOM OOM OOM OOM
CLM 15.68 0.584 0.523 18.7GB 25.1GB 600k 24.2M
Ours 21.59 0.711 0.396 16.4GB 18.0GB 600k 136.7M
MC-smaller-city+ (15.1k images)
HorizonGS 15.28 0.601 0.435 13.9GB 23.6GB 1.2M 34.1M
OctreeGS 15.93 0.560 0.556 26.6GB 98.7GB 310k N/A
Ours 21.71 0.710 0.370 15.7GB 17.2GB 310k 60M
H-3DGS 14.57 0.516 0.563 97.0GB 15.1GB 4M 83.1M
Ours 20.59 0.658 0.468 20.7GB 22.1GB 310k 24.7M
Uni10k (10.1k images)
CityGaussian 19.78 0.642 0.414 3.7GB 14.6GB 300k 2.2M
H-3DGS 19.96 0.705 0.319 39.2GB 26.5GB 255k 32.5M
OctreeGS 20.06 0.602 0.341 25.3GB 22.9GB 160k N/A
CLM-GS 19.89 0.608 0.455 46.7GB 39.7GB 160k 12.6M
HorizonGS 21.72 0.722 0.326 10.3GB 44.3GB 400k N/A
Ours 21.32 0.718 0.336 22.6GB 31.3GB 160k 28.6M
5. Evaluation
Our method is designed to enable seamless training and rendering on ultra-large-scale scenes comprising tens of thousands of views captured at vastly different scales. Unfortunately, most datasets of sufficient size contain either street-level or aerial views—but not both. To address this gap, we captured the campus of Udine University with over 10 000 images at 4k resolution from varying aerial heights and street views in the Uni10k scene. We also introduce MC-small-city+, which expands the small-city scene from the MatrixCity (Li et al., 2023) dataset with new high-altitude aerial views covering the entire scene, resulting in 42.2k images spanning hundreds of buildings. Since some methods run out-of-memory on scenes of this scale, we additionally construct a subset of MC-small-city+ covering about a third of the area (15.1k images, MC-smaller-city+). Additionally, we present results on large-scale street-view and indoor scenes from the Hierarchical 3DGS (Kerbl et al., 2024) and OccluGaussian (Liu et al., 2025) datasets. Results on the aerial datasets UrbanScene3D (Lin et al., 2022) and Mill19 (Turki et al., 2022), as well as the smaller-scale H-3DGS single chunk scene are included in Suppl. A.3 along with additional details on all scenes.
We choose recent divide-and-conquer based 3DGS methods CityGaussian (Liu et al., 2024a), H-3DGS (Kerbl et al., 2024), the large-scale neural Gaussian method OctreeGS (Ren et al., 2024) and out-of-core training method CLM-GS (Zhao et al., 2026) as baselines. Because OccluGaussian (Liu et al., 2025) has not released code yet, we compare only against their self-reported results on their dataset. To enable training on scenes of this scale, it was necessary to modify CityGaussian, HorizonGS and OctreeGS to load images from disk instead of caching them fully in RAM or VRAM. These changes affect training throughput but do not alter optimization behavior or final reconstruction quality. In general, the hyperparameters suggested for large-scale scenes were used for the experiments. Exact details on the training setting and configuration files are included in the supplemental material. Evaluations were run on a single H200 GPU with 141 GB VRAM to allow baselines methods to complete training. Where applicable, we report rendering performance on consumer GPUs to demonstrate practical deployability.
5.1. Results
Qualitative comparisons are shown in Figure 12 and quantitative comparisons in Tables 1 and 2. The outputs of H-3DGS show significant floating and ghosting artifacts. While H-3DGS performs well on individual chunks of MC-smaller-city+, oversized floaters survive the merging procedure and obscure the majority of test images (cf. Figure 7). Floaters and ghosting artifacts can also be found in the results on other scenes. The merging procedure of CityGaussian, which is designed for aerial-only datasets, discards most of the trained Gaussians to avoid chunk artifacts, resulting in mostly artifact-free views and low memory requirements, but noticeably blurry results. A LoD of Gaussians reconstructs novel views for joined street- and aerial data with a high degree of detail while avoiding chunk-based artifacts, resulting in a consistent improvement in quality metrics across scenes. Seamless training also enables faster convergence, substantially reducing the number of training iterations required compared to divide-and-conquer based methods. Results on the OccluGaussian dataset further demonstrate generalizability to indoor environments, which are particularly difficult for divide-and-conquer based methods. All experiments of our method were run on an RTX 3090 GPU, except Uni10k, which required more VRAM for the 4k training images. Lowering the resolution to HD reduces peak VRAM to 20GB for the same parameters.
CLM-GS also employs out-of-core memory during training, but relies solely on frustum culling to limit memory pressure. On Uni10k and MC-small-city+, this results in substantially higher VRAM usage per Gaussian than our method. Further, unlike OctreeGS (Ren et al., 2024) and our approach, their densification is not tailored to large-scale scene reconstruction from sparse views and point clouds. On MatrixCity, CLM-GS instead initializes from a fully dense synthetic point cloud; larger non-aerial scenes (e.g. Campus) fail to reconstruct under all configurations provided with their code.
HorizonGS (Jiang et al., 2025) is not competitive on the MC-city+ tests due to the large scale of the dataset and its diversity of viewpoints. While the authors report results for MatrixCity, only a single block is used in their evaluation. However, it performs competitively on Uni10k, for which it is explicitly designed: the method targets mixed street- and aerial-view scenarios and relies on ground-truth labels to primarily target street-level views for densification. Our universal approach—without additional supervision—closely matches HorizonGS on Uni10k and drastically outperforms it on both MC-city+ scenes. Furthermore, rendering the neural HorizonGS representation is impractical due to a costly preprocessing step: while enabling low VRAM usage, it results in render times of up to 10 seconds per frame.